# Deformation quantization and perverse sheaves

### v1

@@ 3.2. Specialization and microlocalization
As shown in [Kas83], there is a natural embedding
@@
Denote by $\operatorname{gr}_{Z}\mathcal{M}$ the associated graded module with respect to the canonical $V$-filtration $V^{k}_{Z}\mathcal{M}$. We refer to [Kas83, Mon96] for details on the following functors.
@@
on $\tilde{\Lambda}$. By [KK81, Lemma 1.5.1] there is a natural embedding
@@
Consider the Euler vector field on $\mathrm{T}^{*}X$. Since $\tilde{\Lambda}\subset\mathrm{T}^{*}X$ is homogeneous, the Lie derivative along the Euler vector field defines an element $\theta^{\prime}\in\mathcal{D}_{\tilde{\Lambda}}^{K^{1/2}_{\tilde{\Lambda}|X}}$. We denote by $\theta^{\prime}\in\widehat{\mathcal{E}}_{X}(1)|_{\Lambda}$ any preimage (such a lift is characterized by [KK81, Equations (1.5.5) and (1.5.6)]). If $\mathcal{M}$ is a regular holonomic ^$\widehat{\mathcal{E}}_{X}$-module and $u\in\mathcal{M}$ is a local nonvanishing section, then by [KK80, Theorem 4.1.1] there is a polynomial $b_{u}(s)\in\mathbf{C}[s]$ such that $b_{u}(\theta^{\prime})u\in\widehat{\mathcal{E}}_{\Lambda/X}(-1)u$. The order $\mathrm{ord}_{\Lambda}(u)$ of $u\in\mathcal{M}$ is the set of zeros of its $b$-function. For a fixed $\alpha\in\mathbf{R}$ the canonical $V$-filtration of $\mathcal{M}$ is given by the subspaces
@@
Fix $\alpha=\operatorname{codim}Z/2+1$. Then we may consider the $V$-filtration on $\mathcal{M}$ along $Z$ or its microlocal version on $\tilde{\mathcal{M}}$ along $\Lambda$. They are related as follows (see [Mon96, Remark 1.2.4]): the natural map $\pi^{-1}\mathcal{M}\rightarrow\tilde{\mathcal{M}}$ induces an isomorphism
@@ 3.4. Microdifferential modules
We have the following important classification result of microdifferential modules regular along a Legendrian (see [Kas86, Theorem 10.3.1]). Let
@@
be the full substack of twisted $D$-modules on $\tilde{\Lambda}$ which are lisse, i.e. whose underlying $\mathcal{O}$-module is a finite rank vector bundle.
@@
between the stack of ^$\widehat{\mathcal{E}}^{\sqrt{v}}_{X}$-modules regular along $\Lambda$ and lisse twisted $D$-modules on $\tilde{\Lambda}$.
@@ 3.5. Quantized contact transformations
To define the stack of microdifferential modules for a general contact manifold $Y$, Kashiwara in [Kas96] uses the theory of quantized contact transformations developed in [SKK73] that we now recall.
@@
By [Kas03, Theorem 8.11] quantized contact transformations exist locally. We will also need a half-twisted version as follows.
@@
Suppose $V_{X}\subset\mathrm{P}^{*}X$ and $V_{Y}\subset\mathrm{P}^{*}Y$ are open subsets together with a contactomorphism $\psi\colon V_{X}\xrightarrow{\sim}V_{Y}$. A normal-$\ast$-preserving quantized contact transformation over $\psi$ is an isomorphism
@@
Consider a normal-$\ast$-preserving quantized contact transformation $\Psi$ over $\psi\colon V_{X}\xrightarrow{\sim}V_{Y}$ as in the above definition. Suppose $\Lambda_{X}\subset V_{X}$ and $\Lambda_{Y}\subset V_{Y}$ are two Legendrians such that $\Lambda_{Y}=\psi(\Lambda_{X})$. Then a normal-$\ast$-preserving quantized contact transformation $\Psi\colon\widehat{\mathcal{E}}^{\sqrt{v}}_{Y}|_{V_{Y}}\xrightarrow{\sim}\psi_{*}\widehat{\mathcal{E}}^{\sqrt{v}}_{X}|_{V_{X}}$ restricts to an isomorphism
@@
If $\Psi$ is a normal-$\ast$-preserving quantized contact transformation, then
@@
The correspondence between quantized contact transformations and pairs $(\mathcal{M},K)$ is established in [Kas03, Chapter 8.2].
@@
To show that $\sigma_{\Lambda}(K)=\sqrt{\mathrm{vol}_{\mathrm{T}^{*}Y}}$, using the notation of [Kas03, Chapter 8.6] we have to prove that
@@ 3.6. Globalization
Let $Y$ be a contact manifold with a symplectization $\gamma\colon\tilde{Y}\rightarrow Y$. The papers [Kas96, PS04] have defined the algebroid $\mathsf{E}_{Y}$ over $Y$ uniquely determined by the following properties:
@@
    $\mathsf{E}_{Y}|_{U}\cong\widehat{\mathcal{E}}^{\sqrt{v}}_{X}|_{U}.$
@@
Consider a Legendrian submanifold $\Lambda\subset Y$. As the notion of regularity of ^$\widehat{\mathcal{E}}^{\sqrt{v}}_{X}$-modules is local and invariant under quantized contact transformations, it extends to $\mathsf{E}_{Y}$-modules. Moreover, using 3.14 we obtain a global version of the microlocalization functor
$\mu^{dR}_{\Lambda}\colon\mathrm{Mod}_{\mathrm{rh}}(\mathsf{E}_{Y})\longrightarrow\gamma_{*}\mathrm{Mod}_{\mathrm{rh}}\left(\mathcal{D}^{\sqrt{v}}_{\tilde{\Lambda}}\right).$
Using this functor one obtains the following classification result (see [DS07, Corollary 6.4]).
@@
$\mu_{\Lambda}^{dR}\colon\mathrm{Mod}_{\Lambda,\mathrm{rh}}(\mathsf{E}_{Y})\xrightarrow{\sim}\gamma_{*}\mathrm{Mod}_{\mathrm{lisse}}\left(\mathcal{D}^{\sqrt{v}}_{\tilde{\Lambda}}\right)$
between the stack of $\mathsf{E}_{Y}$-modules regular along $\Lambda$ and the stack of lisse twisted $D$-modules on $\tilde{\Lambda}$.
The antiinvolution $\ast$ on ^$\widehat{\mathcal{E}}^{\sqrt{v}}_{X}$ globalizes to an isomorphism of algebroids $\mathsf{E}_{Y}\cong\mathsf{E}_{Y}^{\mathrm{op}}$ for any contact manifold $Y$. Thus, any left $\mathsf{E}_{Y}$-module $\mathsf{M}$ can be considered as a right $\mathsf{E}_{Y}$-module and vice versa.
@@ 4.1. WKB-differential operators
of $\mathbf{C}(\!(\hbar)\!)$-algebras of formal WKB-differential operators introduced in [PS04], where $\rho\colon\mathrm{J}^{1}X\rightarrow\mathrm{T}^{*}X$ and $\hbar=\partial_{t}^{-1}$.
@@
We have that $\rho_{*}(\widehat{\mathcal{E}}_{X\times\mathbf{C}}(0))$ is $\hbar$-complete and has no $\hbar$-torsion. Therefore, by [KS12, Corollary 1.5.7] it is cohomologically $\hbar$-complete. By [KS12, Theorem 1.6.6] in this case it is enough to prove flatness modulo $\hbar$, i.e. it is enough to prove that $\rho_{*}\mathcal{O}_{\mathrm{J}^{1}X}$ is flat as a $\mathcal{O}_{\mathrm{T}^{*}X}$-module. Since the map $\mathrm{J}^{1}X\rightarrow\mathrm{T}^{*}X$ is open, it is flat by [BS76, Theorem V.2.13]. ∎
@@
Let us now explain how to glue these sheaves of algebras. Suppose $U_{X}\subset\mathrm{T}^{*}X$ and $U_{Y}\subset\mathrm{T}^{*}Y$ are open subsets together with a symplectomorphism $\phi\colon U_{X}\xrightarrow{\sim}U_{Y}$. Suppose the symplectomorphism $\phi$ lifts to an isomorphism of contactifications $\psi\colon V_{X}=\rho^{-1}(U_{X})\xrightarrow{\sim}V_{Y}=\rho^{-1}(U_{Y})$, where $\rho\colon\mathrm{J}^{1}X\rightarrow\mathrm{T}^{*}X$ and $\rho\colon\mathrm{J}^{1}Y\rightarrow\mathrm{T}^{*}Y$ are the standard contactifications. Recall that in this case $\psi$ is compatible with the Hamiltonian $\mathbf{C}$-actions. The following is [PS04, Lemma 8.5].
@@
Consider an isomorphism of contactifications $\psi\colon V_{X}=\rho^{-1}(U_{X})\xrightarrow{\sim}V_{Y}=\rho^{-1}(U_{Y})$ as above. Then there is a normal-$\ast$-preserving quantized contact transformation
@@
Consider the setting of 4.5. Then we may restrict the $\ast$-preserving quantized contact transformation $\Psi$ to a normal-$\ast$-preserving quantized symplectic transformation
@@ 4.2. Globalization
Let $S$ be a symplectic manifold. The authors of [PS04] have constructed a $\mathbf{C}(\!(\hbar)\!)$-algebroid $\mathsf{W}_{S}$ on $S$ with the following properties:
@@
    $\mathsf{W}_{S}|_{U}\cong\widehat{\mathcal{W}}^{\sqrt{v}}_{X}|_{U}.$
@@
    $\mathsf{W}_{S}\subset\rho_{*}\mathsf{E}_{Y}.$
The antiinvolution $\ast$ on $\widehat{\mathcal{W}}^{\sqrt{v}}_{X}$ globalizes to a $\mathbf{C}$-linear isomorphism of algebroids $\mathsf{W}_{S}\cong\mathsf{W}^{\mathrm{op}}_{S}$ which sends $\hbar\mapsto-\hbar$. As for $\mathsf{E}_{Y}$-modules, we freely pass between left and right $\mathsf{W}_{S}$-modules using this antiinvolution.
@@
A $\mathsf{W}_{S}$-module is holonomic if its support is a Lagrangian subvariety of $S$.
For a Lagrangian subvariety $L\subset S$ we denote by $\mathsf{W}_{L/S}\subset\mathsf{W}_{S}|_{L}$ generated by
$I_{L}=\{P\in\mathsf{W}_{S}(1)|_{L}\mid\sigma_{1}(P)|_{L}=0\}.$
@@
A holonomic $\mathsf{W}_{S}$-module $\mathcal{M}$ is regular holonomic if locally there is a coherent $\mathsf{W}_{S}(0)$-module $\mathcal{M}_{0}\subset\mathcal{M}$ and a Lagrangian subvariety $L\subset S$ containing $\operatorname{supp}\mathcal{M}$ such that $I_{L}\mathcal{M}_{0}\subset\mathcal{M}_{0}$ and $\mathsf{W}_{S}\mathcal{M}_{0}=\mathcal{M}$.
@@
Let $L\subset S$ be a Lagrangian subvariety and $\Lambda\subset Y$ be its contactification with $\rho\colon Y\rightarrow S$ the projection map. If $\mathcal{M}$ is a $\mathsf{E}_{Y}$-module supported on $\Lambda$, then $\mathcal{M}$, regarded as a $\mathsf{W}_{S}$-module via the embedding $\mathsf{W}_{S}\subset\rho_{*}\mathsf{E}_{Y}$, is supported on $L$ and is, therefore, holonomic. This gives a forgetful functor
$\rho_{*}\mathrm{Mod}_{hol}(\mathsf{E}_{Y})\longrightarrow\mathrm{Mod}_{hol}(\mathsf{W}_{S})$
from holonomic $\mathsf{E}_{Y}$-modules to holonomic $\mathsf{W}_{S}$-modules. Since $\mathsf{W}_{L/S}\subset\rho_{*}\mathsf{E}_{\Lambda/Y}$, this forgetful functor restricts to a functor
$\rho_{*}\mathrm{Mod}_{\mathrm{rh}}(\mathsf{E}_{Y})\longrightarrow\mathrm{Mod}_{\mathrm{rh}}(\mathsf{W}_{S}).$
The authors of [DK11] have introduced the stack $\mathrm{RH}_{S}$ of regular holonomic modules on $S$ as follows:
  • Let $L\subset S$ be a Lagrangian subvariety and $\Lambda\subset Y$ its contactification with projection $\rho\colon Y\rightarrow S$. We define
    $\mathrm{RH}_{S,L}=\rho_{*}\mathrm{Mod}_{\Lambda,\mathrm{rh}}(\mathsf{E}_{Y})=\mathrm{Mod}_{\Lambda,\mathrm{rh}}(\rho_{*}\mathsf{E}_{Y}),$
    the stack of regular holonomic $\mathsf{E}_{Y}$-modules supported on $\Lambda$.
  • As the contactification of $L\subset S$ exists locally and is unique by 1.12, the stack $\mathrm{RH}_{S,L}$ only depends on $L\subset S$.
  • Lagrangian subvarieties of $S$ form a filtered system and we define
    $\mathrm{RH}_{S}=\operatorname*{colim}_{L}\mathrm{RH}_{S,L}.$
Example 4.10.
Let $L\subset S$ be a Lagrangian submanifold. The stack $\mathrm{RH}_{S,L}$ consists of $\mathsf{E}_{Y}$-modules regular along a contactification $\Lambda\subset Y$ of $L\subset S$. By 3.16 microlocalization establishes an equivalence
$\mathrm{RH}_{S,L}\cong\gamma_{*}\mathrm{Mod}_{\mathrm{lisse}}(\mathcal{D}^{\sqrt{v}}_{L\times\mathbf{C}^{\times}}),$
where $\gamma\colon L\times\mathbf{C}^{\times}\rightarrow L$ is the projection on the first factor. Half-twisted $D$-modules along $\mathbf{C}^{\times}$ whose underlying $\mathcal{O}$-module is a line bundle are classified by their order $\lambda\in\mathbf{C}/\mathbf{Z}$. In particular, if $L$ is equipped with an orientation data and given $\lambda\in\mathbf{C}/\mathbf{Z}$, there is a canonical module
$\mathsf{M}^{\lambda}_{L}\in\mathrm{RH}_{S,L}\subset\mathrm{RH}_{S}$
simple along $L$ with order $\lambda$.

### v2

@@ 3.2. Specialization and microlocalization
As shown in [KashiwaraVanishing], there is a natural embedding
@@
Denote by $\operatorname{gr}_{Z}\mathcal{M}$ the associated graded module with respect to the canonical $V$-filtration $V^{k}_{Z}\mathcal{M}$. We refer to [KashiwaraVanishing, MonteiroFernandes] for details on the following functors.
@@
on $\tilde{\Lambda}$. By [KashiwaraKawaiHolonomic, Lemma 1.5.1] there is a natural embedding
@@
Consider the Euler vector field on $\mathrm{T}^{*}X$. Since $\tilde{\Lambda}\subset\mathrm{T}^{*}X$ is homogeneous, the Lie derivative along the Euler vector field defines an element $\theta^{\prime}\in\mathcal{D}_{\tilde{\Lambda}}^{K^{1/2}_{\tilde{\Lambda}|X}}$. We denote by $\theta^{\prime}\in\widehat{\mathcal{E}}_{X}(1)|_{\Lambda}$ any preimage (such a lift is characterized by [KashiwaraKawaiHolonomic, Equations (1.5.5) and (1.5.6)]). If $\mathcal{M}$ is a regular holonomic ^$\widehat{\mathcal{E}}_{X}$-module and $u\in\mathcal{M}$ is a local nonvanishing section, then by [KashiwaraKawaiMicrolocalization, Theorem 4.1.1] there is a polynomial $b_{u}(s)\in\mathbf{C}[s]$ such that $b_{u}(\theta^{\prime})u\in\widehat{\mathcal{E}}_{\Lambda/X}(-1)u$. The order $\mathrm{ord}_{\Lambda}(u)$ of $u\in\mathcal{M}$ is the set of zeros of its $b$-function. For a fixed $\alpha\in\mathbf{R}$ the canonical $V$-filtration of $\mathcal{M}$ is given by the subspaces
@@
Fix $\alpha=\operatorname{codim}Z/2+1$. Then we may consider the $V$-filtration on $\mathcal{M}$ along $Z$ or its microlocal version on $\tilde{\mathcal{M}}$ along $\Lambda$. They are related as follows (see [MonteiroFernandes, Remark 1.2.4]): the natural map $\pi^{-1}\mathcal{M}\rightarrow\tilde{\mathcal{M}}$ induces an isomorphism
@@ 3.4. Microdifferential modules
We have the following important classification result of microdifferential modules regular along a Legendrian. Let
@@
be the full substack of twisted $D$-modules on $\tilde{\Lambda}$ which are local systems, i.e. whose underlying $\mathcal{O}$-module is a finite rank vector bundle. The following statement combines [KashiwaraIntroduction, Theorem 10.3.1] and [KashiwaraKawaiMicrolocal, Theorem 7.9].
@@
between the stack of ^$\widehat{\mathcal{E}}^{\sqrt{v}}_{X}$-modules regular along $\Lambda$ and twisted local systems on $\tilde{\Lambda}$. Moreover, if $\mathcal{M}$ is a regular holonomic ^$\widehat{\mathcal{E}}^{\sqrt{v}}_{X}$-module and $\mathcal{L}$ a ^$\widehat{\mathcal{E}}^{\sqrt{v}}_{X}$-module regular along $\Lambda$, then microlocalization defines an isomorphism
$\mathbb{R}\mathcal{H}om_{\widehat{\mathcal{E}}^{\sqrt{v}}_{X}}(\mathcal{M},\mathcal{L})\xrightarrow{\sim}\gamma_{*}\mathbb{R}\mathcal{H}om_{\mathcal{D}^{\sqrt{v}}_{\tilde{\Lambda}}}(\mu_{\Lambda}^{dR}(\mathcal{M}),\mu_{\Lambda}^{dR}(\mathcal{L})).$
@@ 3.5. Quantized contact transformations
To define the stack of microdifferential modules for a general contact manifold $Y$, Kashiwara in [KashiwaraContact] uses the theory of quantized contact transformations developed in [SatoKawaiKashiwara] that we now recall.
@@
By [KashiwaraDmodules, Theorem 8.11] quantized contact transformations exist locally. We will also need a half-twisted version as follows.
@@
Suppose $V_{X}\subset\mathrm{P}^{*}X$ and $V_{Y}\subset\mathrm{P}^{*}Y$ are open subsets together with a contactomorphism $\psi\colon V_{X}\xrightarrow{\sim}V_{Y}$. A $\ast$-preserving quantized contact transformation over $\psi$ is an isomorphism
@@
Consider a $\ast$-preserving quantized contact transformation $\Psi$ over $\psi\colon V_{X}\xrightarrow{\sim}V_{Y}$ as in the above definition. Suppose $\Lambda_{X}\subset V_{X}$ and $\Lambda_{Y}\subset V_{Y}$ are two Legendrians such that $\Lambda_{Y}=\psi(\Lambda_{X})$. Then a $\ast$-preserving quantized contact transformation $\Psi\colon\widehat{\mathcal{E}}^{\sqrt{v}}_{Y}|_{V_{Y}}\xrightarrow{\sim}\psi_{*}\widehat{\mathcal{E}}^{\sqrt{v}}_{X}|_{V_{X}}$ restricts to an isomorphism
@@
If $\Psi$ is a $\ast$-preserving quantized contact transformation, then
@@
The correspondence between quantized contact transformations and pairs $(\mathcal{M},K)$ is established in [KashiwaraDmodules, Chapter 8.2].
@@
To show that $\sigma_{\Lambda}(K)=\sqrt{\mathrm{vol}_{\mathrm{T}^{*}Y}}$, using the notation of [KashiwaraDmodules, Chapter 8.6] we have to prove that
@@ 3.6. Globalization
Let $Y$ be a contact manifold with a symplectization $\gamma\colon\tilde{Y}\rightarrow Y$. The papers [KashiwaraContact, PoleselloSchapira] have defined the algebroid $\widehat{\mathsf{E}}_{Y}$ over $Y$ uniquely determined by the following properties:
@@
    $\widehat{\mathsf{E}}_{Y}|_{U}\cong\widehat{\mathcal{E}}^{\sqrt{v}}_{X}|_{U}.$
@@
Consider a Legendrian submanifold $\Lambda\subset Y$. As the notion of regularity of ^$\widehat{\mathcal{E}}^{\sqrt{v}}_{X}$-modules is local and invariant under quantized contact transformations, it extends to $\widehat{\mathsf{E}}_{Y}$-modules. Moreover, using 3.14 we obtain a global version of the microlocalization functor
$\mu^{dR}_{\Lambda}\colon\mathrm{Mod}_{\mathrm{rh}}(\widehat{\mathsf{E}}_{Y})\longrightarrow\gamma_{*}\mathrm{Mod}_{\mathrm{rh}}\left(\mathcal{D}^{\sqrt{v}}_{\tilde{\Lambda}}\right).$
Using this functor one obtains the following classification result generalizing 3.10 (see [DAgnoloSchapira, Corollary 6.4]).
@@
$\mu_{\Lambda}^{dR}\colon\mathrm{Mod}_{\Lambda,\mathrm{rh}}(\widehat{\mathsf{E}}_{Y})\xrightarrow{\sim}\gamma_{*}\mathrm{Mod}_{\mathrm{lisse}}\left(\mathcal{D}^{\sqrt{v}}_{\tilde{\Lambda}}\right)$
between the stack of $\widehat{\mathsf{E}}_{Y}$-modules regular along $\Lambda$ and the stack of twisted local systems on $\tilde{\Lambda}$. Moreover, if $\mathcal{M}$ is a regular holonomic $\widehat{\mathsf{E}}_{Y}$-module and $\mathcal{L}$ is an $\widehat{\mathsf{E}}_{Y}$-module regular along $\Lambda$, then microlocalization defines an isomorphism
$\mathbb{R}\mathcal{H}om_{\widehat{\mathsf{E}}_{Y}}(\mathcal{M},\mathcal{L})\xrightarrow{\sim}\gamma_{*}\mathbb{R}\mathcal{H}om_{\mathcal{D}^{\sqrt{v}}_{\tilde{\Lambda}}}(\mu_{\Lambda}^{dR}(\mathcal{M}),\mu_{\Lambda}^{dR}(\mathcal{L})).$
The antiinvolution $\ast$ on ^$\widehat{\mathcal{E}}^{\sqrt{v}}_{X}$ globalizes to an isomorphism of algebroids $\widehat{\mathsf{E}}_{Y}\cong\widehat{\mathsf{E}}_{Y}^{\mathrm{op}}$ for any contact manifold $Y$. Thus, any left $\widehat{\mathsf{E}}_{Y}$-module $\mathsf{M}$ can be considered as a right $\widehat{\mathsf{E}}_{Y}$-module and vice versa.
@@ 4.1. WKB-differential operators
of $\mathbf{C}(\!(\hbar)\!)$-algebras of formal WKB-differential operators introduced in [PoleselloSchapira], where $\rho\colon\mathrm{J}^{1}X\rightarrow\mathrm{T}^{*}X$ and $\hbar=\partial_{t}^{-1}$.
@@
We have that $\rho_{*}(\widehat{\mathcal{E}}_{X\times\mathbf{C}}(0))$ is $\hbar$-complete and has no $\hbar$-torsion. Therefore, by [KashiwaraSchapira, Corollary 1.5.7] it is cohomologically $\hbar$-complete. By [KashiwaraSchapira, Theorem 1.6.6] in this case it is enough to prove flatness modulo $\hbar$, i.e. it is enough to prove that $\rho_{*}\mathcal{O}_{\mathrm{J}^{1}X}$ is flat as a $\mathcal{O}_{\mathrm{T}^{*}X}$-module. Since the map $\mathrm{J}^{1}X\rightarrow\mathrm{T}^{*}X$ is open, it is flat by [BanicaStanasila, Theorem V.2.13]. ∎
@@
Let us now explain how to glue these sheaves of algebras. Suppose $U_{X}\subset\mathrm{T}^{*}X$ and $U_{Y}\subset\mathrm{T}^{*}Y$ are open subsets together with a symplectomorphism $\phi\colon U_{X}\xrightarrow{\sim}U_{Y}$. Suppose the symplectomorphism $\phi$ lifts to an isomorphism of contactifications $\psi\colon V_{X}=\rho^{-1}(U_{X})\xrightarrow{\sim}V_{Y}=\rho^{-1}(U_{Y})$, where $\rho\colon\mathrm{J}^{1}X\rightarrow\mathrm{T}^{*}X$ and $\rho\colon\mathrm{J}^{1}Y\rightarrow\mathrm{T}^{*}Y$ are the standard contactifications. Recall that in this case $\psi$ is compatible with the Hamiltonian $\mathbf{C}$-actions. The following is [PoleselloSchapira, Lemma 8.5].
@@
Consider an isomorphism of contactifications $\psi\colon V_{X}=\rho^{-1}(U_{X})\xrightarrow{\sim}V_{Y}=\rho^{-1}(U_{Y})$ as above. Then there is a $\ast$-preserving quantized contact transformation
@@
Consider the setting of 4.5. Then we may restrict the $\ast$-preserving quantized contact transformation $\Psi$ to a $\ast$-preserving quantized symplectic transformation
@@ 4.2. Globalization
Let $S$ be a symplectic manifold. The authors of [PoleselloSchapira] have constructed a $\mathbf{C}(\!(\hbar)\!)$-algebroid $\widehat{\mathsf{W}}_{S}$ on $S$ with the following properties:
@@
    $\widehat{\mathsf{W}}_{S}|_{U}\cong\widehat{\mathcal{W}}^{\sqrt{v}}_{X}|_{U}.$
@@
    $\widehat{\mathsf{W}}_{S}\subset\rho_{*}\widehat{\mathsf{E}}_{Y}.$
The antiinvolution $\ast$ on $\widehat{\mathcal{W}}^{\sqrt{v}}_{X}$ globalizes to a $\mathbf{C}$-linear isomorphism of algebroids $\widehat{\mathsf{W}}_{S}\cong\widehat{\mathsf{W}}^{\mathrm{op}}_{S}$ which sends $\hbar\mapsto-\hbar$. As for $\widehat{\mathsf{E}}_{Y}$-modules, we freely pass between left and right $\widehat{\mathsf{W}}_{S}$-modules using this antiinvolution.
@@
A $\widehat{\mathsf{W}}_{S}$-module is holonomic if its support is a Lagrangian subvariety of $S$.
For a Lagrangian subvariety $L\subset S$ we denote by $\widehat{\mathsf{W}}_{L/S}\subset\widehat{\mathsf{W}}_{S}|_{L}$ generated by
$I_{L}=\{P\in\widehat{\mathsf{W}}_{S}(1)|_{L}\mid\sigma_{1}(P)|_{L}=0\}.$
@@
A holonomic $\widehat{\mathsf{W}}_{S}$-module $\mathcal{M}$ is regular holonomic if locally there is a coherent $\widehat{\mathsf{W}}_{S}(0)$-module $\mathcal{M}_{0}\subset\mathcal{M}$ and a Lagrangian subvariety $L\subset S$ containing $\operatorname{supp}\mathcal{M}$ such that $I_{L}\mathcal{M}_{0}\subset\mathcal{M}_{0}$ and $\widehat{\mathsf{W}}_{S}\mathcal{M}_{0}=\mathcal{M}$.
@@
Let $L\subset S$ be a Lagrangian subvariety and $\Lambda\subset Y$ be its contactification with $\rho\colon Y\rightarrow S$ the projection map. If $\mathcal{M}$ is a $\widehat{\mathsf{E}}_{Y}$-module supported on $\Lambda$, then $\mathcal{M}$, regarded as a $\widehat{\mathsf{W}}_{S}$-module via the embedding $\widehat{\mathsf{W}}_{S}\subset\rho_{*}\widehat{\mathsf{E}}_{Y}$, is supported on $L$ and is, therefore, holonomic. This gives a forgetful functor
$\rho_{*}\mathrm{Mod}_{hol}(\widehat{\mathsf{E}}_{Y})\longrightarrow\mathrm{Mod}_{hol}(\widehat{\mathsf{W}}_{S})$
from holonomic $\widehat{\mathsf{E}}_{Y}$-modules to holonomic $\widehat{\mathsf{W}}_{S}$-modules. Since $\widehat{\mathsf{W}}_{L/S}\subset\rho_{*}\widehat{\mathsf{E}}_{\Lambda/Y}$, this forgetful functor restricts to a functor
$\rho_{*}\mathrm{Mod}_{\mathrm{rh}}(\widehat{\mathsf{E}}_{Y})\longrightarrow\mathrm{Mod}_{\mathrm{rh}}(\widehat{\mathsf{W}}_{S}).$